# Lattice Hamiltonians and Stray Interactions Within Quantum Processors

### v2

@@ V $ZZZ$ Superiority
This analysis highlights the intricate balance and interplay of multi-qubit interactions in quantum circuits. It underscores the importance of considering both two-body and three-body interactions in circuit design, especially in scenarios aiming for high-fidelity gate operations. The findings suggest that a more nuanced understanding of these interactions can lead to optimized strategies for quantum computing, avoiding oversimplified assumptions of interaction hierarchies and paving the way for more effective quantum gate designs.
VI Conclusion
@@
In conclusion, this study highlights the critical importance of considering the lattice Hamiltonian and many-body couplings in the design of quantum circuits. This allows for fine-tuning circuit parameters to mitigate the determinal effcts of stray couplings. More specifically we showed that both three-body stray couplings can be as important as two-body interactions and it can effectively control strategic adjustment of circuit for better gate performance, paving the way for the implementation of high-fidelity two-qubit gates.
@@
Furthermore, we have shown that the dominance of three-body $ZZZ$ interactions can be realized within the quasi-dispersive regime, suggesting the possibility of achieving optimal quantum performance even outside the near-degenerate regimes. These findings offer significant contributions to the fields of quantum system design and algorithm optimization, promising enhancements in the efficiency and robustness of quantum computing technologies.
@@ Acknowledgement
We appreciate our constructive discussions with John Preskill, Roger Melko, and Michele Mosca.   M.H. Ansari would like to express gratitude for the technical support and enriching discussions with Joséphine Pazem that greatly contributed to the development of this work. X. Xu was partially supported by the Project GeQCoS, Grant No. 13N15685. Manabputra’s involvement was partially supported by the Kishore Vaigyanik Protsahan Yojana (KVPY) Fellowship, sponsored by the Department of Science & Technology, Government of India. This research received funding from Horizon Europe (HORIZON) Project: 101113946 OpenSuperQPlus100.
@@ Appendix B Reduced circuit Hamiltonian
Appendix C Divergence in the three-qubit circuit
Table 1 summarizes all the three-qubit resonances in the circuit in Fig. 3 as observed in the plot in Fig. 4. Presence of a third qubit introduces many more high level divergences that are not captured by the two qubit analysis.
States $|Q_{1}Q_{2}Q_{3}\rangle$ Condition $\omega_{2}/2\pi$ (GHz)
$|001\rangle\sim|010\rangle$ $\Delta_{23}=0$ $5.430$
$|101\rangle\sim|110\rangle$
$|001\rangle\sim|100\rangle$ $\Delta_{13}=0$
$|011\rangle\sim|110\rangle$
$|010\rangle\sim|100\rangle$ $\Delta_{12}=0$ $4.800$
$|011\rangle\sim|101\rangle$
$|011\rangle\sim|020\rangle$ $\Delta_{23}=-\delta_{2}$ $5.430$
$|111\rangle\sim|120\rangle$
$|011\rangle\sim|002\rangle$ $\Delta_{23}=\delta_{3}$ $4.770$
$|111\rangle\sim|102\rangle$
$|011\rangle\sim|200\rangle$ $\Delta_{12}+\Delta_{13}=-\delta_{1}$ $4.170$
$|101\rangle\sim|200\rangle$ $\Delta_{13}=-\delta_{1}$
$|111\rangle\sim|210\rangle$
$|101\rangle\sim|002\rangle$ $\Delta_{13}=\delta_{3}$
$|111\rangle\sim|012\rangle$
$|101\rangle\sim|020\rangle$ $\Delta_{12}-\Delta_{23}=\delta_{2}$ $5.115$
$|110\rangle\sim|020\rangle$ $\Delta_{12}=\delta_{2}$ $5.130$
$|111\rangle\sim|021\rangle$
$|110\rangle\sim|200\rangle$ $\Delta_{12}=-\delta_{2}$ $4.470$
$|111\rangle\sim|201\rangle$
$|110\rangle\sim|002\rangle$ $\Delta_{13}+\Delta_{23}=\delta_{3}$ $5.070$
$|111\rangle\sim|003\rangle$ $\Delta_{13}+\Delta_{23}=3\delta_{3}$ $4.410$
$|111\rangle\sim|030\rangle$ $\Delta_{12}-\Delta_{23}=3\delta_{2}$ $5.445$
$|111\rangle\sim|300\rangle$ $\Delta_{12}+\Delta_{13}=-3\delta_{1}$ $3.510$
Table 1: Summary of three-qubit resonances in the three-qubit subspace. The circuit parameters are same as in Fig. 4.
Appendix D Qubit coupling symmetry
@@
Figure 10 and 11 depict the numerical simulation of Eq. (14) using parameters similar to those in Fig. 4 and Fig. 7(c), respectively. These results exhibit significant consistency, affirming the validity of the identity.
Figure 10: Numerical simulation of $J$-relations in Eq. (14). The circuit parameters are same as those in Fig.  4.
Figure 11: Numerical simulation of $J$-relations in Eq. (14). The circuit parameters are same as those in Fig.  7(c).
Appendix E ZIZ Discrepancy
@@
Appendix F Effective Coupling Discrepancy
@@
In Fig. 14, we describe the landscape with C12 frequency at decoupling points at B and E. In Fig. 15, we describe the landscape with C12 frequency at decoupling points at C and F. Because of the higher order $J^{01}$ and $J^{10}$ interactions, the landscapes have significant features. In the case of the $ZZZ$ interaction, especially when the qubits are strongly coupled, it can go upwards of thousand kilohertz as $\omega_{c23}$ varies, even when $J_{12}$ is zero. $IZZ$ also varies significantly with $\omega_{c23}$ but $ZIZ$ have little dependence on it. In the second row of graphs, when the direct coupling $g_{12}$ is stronger, the zeroness of $ZZI$ disappears and becomes narrower for $ZIZ$. At the decoupling points, $ZZZ$, $ZZI$ and $ZIZ$ are all smaller than 50 kHz whereas $IZZ$ is significantly stronger.
Figure 14: Stray couplings in the triangular circuit with all-to-all connectivity as a function of the frequencies of the resonators C13 and C23 $\omega_{c13}$ and $\omega_{c23}$ respectively. The first row is at $g_{12}/2\pi=4$ MHz with nearly zero $J_{12}$ and $J_{13}$ at spot B: $(\omega_{c12}/2\pi,\omega_{c23}/2\pi,\omega_{c13}/2\pi)=(6.67,5.405,6.15)$ GHz. The second row is at $g_{12}/2\pi=13$ MHz with nearly zero $J_{12}$ and $J_{13}$ at spot E: $(5.443,5.472,6.076)$ GHz. Other circuit parameters are the same as Fig. 4 with $\omega_{2}/2\pi=5.0$ GHz.
Figure 15: Stray couplings in the triangular circuit with all-to-all connectivity as a function of the frequencies of the resonators C23 and C13 i.e. $\omega_{c23}$ and $\omega_{c13}$ respectively. The first row is at $g_{12}/2\pi=4$ MHz with nearly zero $J_{12}$ and $J_{13}$ at spot C: $(\omega_{c12}/2\pi,\omega_{c23}/2\pi,\omega_{c13}/2\pi)=(6.68,5.275,6.154)$ GHz. The second row is at $g_{12}/2\pi=13$ MHz with nearly zero $J_{12}$ and $J_{13}$ at spot F: $(5.446,5.621,6.098)$ GHz. Other circuit parameters are the same as Fig. 4 with $\omega_{2}/2\pi=5.0$ GHz.
@@
Appendix G Three qubits and one coupler
We extend our investigation beyond the case of three-qubit three-coupler architecture. In this section, we study the case of three qubits sharing a coupler, which resembles the one recently studied in Ref. [18]. For simplicity, we consider three transmon qubits coupled to a harmonic oscillator as the shared coupler, as depicted in Fig. 17. Our study aims to examine the behaviour of such a system in response to changes in coupler frequency $\omega_{c}$ and direct coupling, i.e. $g_{13}$. We present the findings in Fig. 18.
Figure 17: Triangle geometry of three qubit interaction. All three qubits are indirectly coupled via one shared coupler, and each pair is directly coupled via capacitors.
(a)(b)(c)(d)(e)
Figure 18: Stray couplings versus coupler C frequency of circuit depicted in Fig. 17. (a) $ZZZ$, (b) $ZZI$, (c) $ZIZ$, (d) $IZZ$, (e) $|ZZZ|/{|ZZ|_{\max}}$. Circuit parameters are: $\omega_{1}/2\pi=4.8$ GHz, $\omega_{2}/2\pi=5.0$ GHz, $\omega_{3}/2\pi=5.1$ GHz, $\delta_{1}/2\pi=\delta_{2}/2\pi=\delta_{3}/2\pi=-330$ MHz with qubit-coupler strength $g_{1c1}/2\pi=g_{1c_{3}}/2\pi=g_{2c_{1}}/2\pi=g_{3c_{3}}/2\pi=85$ MHz and $g_{2c_{2}}/2\pi=g_{3c_{2}}/2\pi=102$ MHz, direct qubit-qubit coupling $g_{12}=g_{23}/2\pi=4$ MHz.

### v3

@@ V $ZZZ$ Superiority
This analysis highlights the intricate balance of multi-qubit interactions in quantum circuits, emphasizing the need to consider both two-body and three-body interactions in circuit design. Especially in scenarios aiming for high-fidelity gate operations, the findings suggest that a nuanced understanding of these interactions can lead to optimized strategies for quantum computing. This approach avoids oversimplified assumptions of interaction hierarchies and paves the way for more effective quantum gate designs.
VI Cross Resonance gate on Lattice
Cross resonance ($CR$) gate is a type of interacting entanglement between two qubits, applied locally on one qubit, namely ‘control’, and makes the other qubit, namely ‘target’, undergoes ±$\pm\hat{X}$ operation, with $\pm$ being subject to the state of the control qubit; i.e. the interacting Hamiltonian between is $ZX$. This can be done in the lab by driving the control qubit with a microwave pulse of frequency of the target qubit [50, 51]. Here we study the impact of microwave driving on the strength of stray couplings.
Applying $CR$ gate on qubits makes an impact on the stray couplings between qubits and this has been previously extensively studied in [52, 45, 53, 54], even in Ref. [45, 55, 56] it has been proposed that the microwave activated part of the stray coupling could take place with the opposite sign of the static stray coupling of no $CR$ gate, so that by tuning circuit parameters one can cancel the two and make the so-called $ZZ$-free qubits, which is yet to be verified experimentally. However, detailed impact of $CR$ gate on the lattice Hamilton in many-body form is unknown.
In this section study the impact of $CR$ gate on $ZZ$ and $ZZZ$ for a system of superconducting qubits. The microwave driving Hamiltonian in the lab frame is $H_{\textup{drive}}=\Omega\cos(\omega_{\textup{target}}t)(\hat{a}_{\textup{control}}+\hat{a}^{\dagger}_{\textup{control}})$. We take this part of Hamiltonian into the relevant frame of non-perturbative block diagonalization, and together with the static Hamiltonian in the basis of $|Q_{\textup{control}},Q_{\textup{target}},Q_{\textup{spectator}}\rangle$, final result will end up in the following general form:
$H$ $=$ $\alpha_{ZZI}\frac{ZZI}{4}+\alpha_{ZIZ}\frac{ZIZ}{4}+\alpha_{IZZ}\frac{IZZ}{4}+\alpha_{ZZZ}\frac{ZZZ}{8}$ (12)
$+\alpha_{ZXI}\frac{ZXI}{2}+\alpha_{ZXZ}\frac{ZXZ}{4}$
Here classical crosstalk is assumed to be canceled. The two-qubit version of these Pauli coefficients —$\alpha_{ijk}$ with one of the three indices being identity and the other two from the set $\{I,Z,X,Y\}$— have been analytically determined by perturbative block diagonalization techniques in [45, 47]. Non-perturbative numerical analysis show that the pairwise Pauli coefficients in the leading order of CR gate amplitude scale as $\alpha_{ZZ}(\Omega)=\alpha_{ZZ}(0)+\eta_{2}\Omega^{2}+\eta_{a}\Omega^{a}$ with coupler frequency dependent power $4\leq a\leq 5$, and $\alpha_{ZX}(\Omega)=\mu_{1}\Omega+\mu_{b}\Omega^{b}$ with $b$ being nearly 3. For more discussion on the method of evaluation and domain of parameters see Ref. [54], Fig. 8. It is evident that $ZZZ$ under the microwave drive of a $CR$ gate changes from its static value by the following form, which can be verified perturbatively: $\alpha_{ZZZ}(\Omega)=\alpha_{ZZZ}(0)+\nu_{2}\Omega^{2}+O(3)$ in the leading order, with $\nu_{2}$ depending on the parameters of circuit and can be positive or negative, meaning that the microwave-activated part of the $ZZZ$ can add or reduce error from the circuit lattice Hamiltonian.
Figure 10: Schematic of thirteen Xmon qubits with the bottom six qubits serving as tunable couplers. See Ref. [6] for more details.
We evaluate the Pauli coefficients in the Lattice Hamiltonian of Eq. (12) numerically. In order to compare our simulation results with experiment we study the recent measurement data of a 7-qubit circuit in Ref. [6], as shown in Fig. 10. Qubits interact by cross-resonance gate. In particular, the authors assess spectator errors through three-qubit tomography in Fig. S7(D). In this case, let $Q_{2}$ be the control qubit, $Q_{3}$ the target, and $Q_{4}$ the target spectator qubit, with no coupler between $Q_{2}$ and $Q_{4}$. Here, we employ our non-perturbative approach to simulate the corresponding Pauli coefficients at the driving amplitude $\Omega/2\pi=18$ MHz, illustrated in Fig. 11. In order to match the measured $ZZZ$ and $ZXZ$ data, we assume $Q_{2}$ and $Q_{4}$ are capacitively coupled with a strength of $g_{24d}/2\pi=4.5$ MHz. Markers depict the extracted experimental data, while curves represent results from our numerical method. Our theory effectively predicts the behavior of such qubit chain devices.
Table 1 summarizes several static and microwave-activated resonances —perturbative divergences— caused in the presence of a third qubit in the circuit, as discussed in Ref. [6]. The static (microwave-activated) resonances are highlighted in vertical blue (black) dashed lines in Fig. 4. The introduction of a third qubit leads to numerous higher-level divergences that are not accounted for in the two-qubit analysis.
[…]
$|001\rangle\sim|010\rangle$ $\Delta_{st}=0$ 00
$|101\rangle\sim|110\rangle$
$|001\rangle\sim|100\rangle$ $\Delta_{cs}=0$ 137
$|011\rangle\sim|110\rangle$
$|011\rangle\sim|020\rangle$ $\Delta_{st}=\delta_{2}$ $-218$
$|111\rangle\sim|120\rangle$
$|011\rangle\sim|002\rangle$ $\Delta_{st}=-\delta_{3}$ $213$
$|111\rangle\sim|102\rangle$
$|011\rangle\sim|200\rangle$ $\Delta_{cs}+\Delta_{cs}=-\delta_{1}$ $56$
$|101\rangle\sim|200\rangle$ $\Delta_{cs}=-\delta_{1}$ $-81$
$|111\rangle\sim|210\rangle$
$|101\rangle\sim|002\rangle$ $\Delta_{cs}=\delta_{3}$ $350$
$|111\rangle\sim|012\rangle$
$|101\rangle\sim|020\rangle$ $\Delta_{ct}+\Delta_{st}=\delta_{2}$ $-355$
$|110\rangle\sim|002\rangle$ $\Delta_{cs}-\Delta_{st}=\delta_{3}$ $175$
$|020\rangle\sim|002\rangle$ $2\Delta_{st}=\delta_{2}-\delta_{3}$ $-2.5$
$|200\rangle\sim|002\rangle$ $2\Delta_{st}-2\Delta_{ct}=\delta_{1}-\delta_{3}$ $134.5$
$|012\rangle\sim|120\rangle$ $2\Delta_{st}=\Delta_{ct}+\delta_{2}-\delta_{3}$ $66$
[…]
$|000\rangle|n_{d}+2\rangle\sim|002\rangle|n_{d}\rangle$ $2\Delta_{st}=-\delta_{3}$ $106.5$
$|000\rangle|n_{d}+2\rangle\sim|101\rangle|n_{d}\rangle$ $\Delta_{st}=-\Delta_{ct}$ $-137$
$|300\rangle|n_{d}+1\rangle\sim|202\rangle|n_{d}\rangle$ $2\Delta_{st}=\Delta_{ct}+2\delta_{1}-\delta_{3}$ $-180$
$|003\rangle|n_{d}+1\rangle\sim|400\rangle|n_{d}\rangle$ $3\Delta_{st}=4\Delta_{ct}+6\delta_{1}-3\delta_{3}$ $-40.3$
Table 1: Summary of three-qubit resonances in the three-qubit subspace. The states are denoted in the following order: $|Q_{\textup{control}}Q_{\textup{target}}Q_{\textup{spectator}}\rangle$. The circuit parameters are same as in Fig. 4.
(a)(b)(c)(d)
Figure 11: Interaction terms as varying the spectator-target detuning $\Delta_{st}$ with fixed CR drive amplitude $\Omega/2\pi=18$ MHz when considering parasitic interactions from the target spectator qubit: (a) $ZZI$, (b) $IZZ$, (c) $ZZZ$ and (d) $ZXI$ ($ZXZ$). Markers represent the extracted experimental data in Ref. [6], curves are the numerical simulation results.
In order to further our understanding of the impact of CR gate on the lattice Hamiltonian of three qubits, we use our software CirQubit to generate numerical data for the cases that experiment has not conducted, such as the static stray coupling in absence of $CR$ gate $\Omega=0$, and the case of $CR$ gate amplitude being $\Omega=31$ MHz, which is almost the double of the experimental data taken at $\Omega=16$ MHz. Results have been plotted in Appendix G Fig. 21, which indicates how the Lattice Pauli coefficients of Eq. (12) depends on the frequency detuning between control and target qubits $\Delta_{st}$. Interestingly we see in Fig. 21 in a large domain of $\Delta_{st}$, by increasing the microwave power, the interactions $ZXI$ and $ZXZ$ become more powerful, while $ZZ$ and $ZZZ$ stray couplings are suppressed. The suppression of $ZZ$’s have been predicted earlier in Ref. [45] known as ‘$ZZ$-freedom.’ Here, we extend that result by demonstrating that it is possible to tune the circuit parameters to suppress the three-body stray coupling $ZZZ$.
VII Conclusion
@@
In conclusion, this study underscores the critical importance of considering the lattice Hamiltonian and many-body couplings in the design and operation of quantum circuits. This approach allows for fine-tuning circuit parameters to mitigate the detrimental effects of stray couplings. Furthermore, we demonstrate that three-body stray couplings can be as significant as two-body interactions, particularly beyond the dispersive regime. Effectively controlling these interactions through strategic adjustments can enhance gate performance and holds potential for the implementation of high-fidelity two-qubit gates.
@@
Additionally, we have shown that the dominance of three-body $ZZZ$ interactions can be realized within the quasi-dispersive regime, suggesting the possibility of achieving optimal quantum performance even outside the near-degenerate regimes. These findings offer significant contributions to the fields of quantum system design and algorithm optimization, promising enhancements in the efficiency and robustness of quantum computing technologies.
@@ Acknowledgement
M Ansari would like to express gratitude for the technical support and enriching discussions with Joséphine Pazem that greatly contributed to the development of this work. X. Xu was partially supported by the Project GeQCoS, Grant No. 13N15685. Manabputra’s involvement was partially supported by the Kishore Vaigyanik Protsahan Yojana (KVPY) Fellowship, sponsored by the Department of Science & Technology, Government of India. This research received funding from Horizon Europe (HORIZON) Project: 101113946 OpenSuperQPlus100.
@@ Appendix B Reduced circuit Hamiltonian
Appendix C Qubit coupling symmetry
@@
Figure 14: Numerical simulation of $J$-relations in Eq. (11). The circuit parameters are same as those in Fig. 4.
Figure 15: Numerical simulation of $J$-relations in Eq. (11). The circuit parameters are same as those in Fig. 9(c).
Figure 14 and 15 depict the numerical simulation of Eq. (11) using parameters similar to those in Fig. 4 and Fig. 9(c), respectively. These results exhibit significant consistency, affirming the validity of the identity.
Appendix D ZIZ Discrepancy
@@
Appendix E Effective Coupling Discrepancy
@@
In Fig. 18(a), we describe the landscape with C12 frequency at decoupling points at B and E. In Fig. 18(b), we describe the landscape with C12 frequency at decoupling points at C and F. Because of the higher order $J^{01}$ and $J^{10}$ interactions, the landscapes have significant features. In the case of the $ZZZ$ interaction, especially when the qubits are strongly coupled, it can go upwards of thousand kilohertz as $\omega_{c23}$ varies, even when $J_{12}$ is zero. $IZZ$ also varies significantly with $\omega_{c23}$ but $ZIZ$ have little dependence on it. In the second row of graphs, when the direct coupling $g_{12}$ is stronger, the zeroness of $ZZI$ disappears and becomes narrower for $ZIZ$. At the decoupling points, $ZZZ$, $ZZI$ and $ZIZ$ are all smaller than 50 kHz whereas $IZZ$ is significantly stronger.
Refer to caption(b)
Figure 18: (a) Stray couplings in the triangular circuit with all-to-all connectivity as a function of the frequencies of the resonators C13 and C23 $\omega_{c13}$ and $\omega_{c23}$ respectively. The first row is at $g_{12}/2\pi=4$ MHz with nearly zero $J_{12}$ and $J_{13}$ at spot B: $(\omega_{c12}/2\pi,\omega_{c23}/2\pi,\omega_{c13}/2\pi)=(6.67,5.405,6.15)$ GHz. The second row is at $g_{12}/2\pi=13$ MHz with nearly zero $J_{12}$ and $J_{13}$ at spot E: $(5.443,5.472,6.076)$ GHz. (b) The third (fourth) row is at $g_{12}/2\pi=4$ (13) MHz with nearly zero $J_{12}$ and $J_{13}$ at spot C: $(\omega_{c12}/2\pi,\omega_{c23}/2\pi,\omega_{c13}/2\pi)=(6.68,5.275,6.154)$ GHz. In the fourth row $J_{12}$ and $J_{13}$ are nearly zero at spot F: $(5.446,5.621,6.098)$ GHz. Other circuit parameters are the same as Fig. 4 with $\omega_{2}/2\pi=5.0$ GHz.
@@
Appendix F Three qubits and one coupler
We extend our investigation beyond the case of three-qubit three-coupler architecture. In this section, we study the case of three qubits sharing a coupler, which resembles the one recently studied in Ref. [27]. For simplicity, we consider three transmon qubits coupled to a harmonic oscillator as the shared coupler, as depicted in Fig. 20(a). Our study aims to examine the behavior of such a system in response to changes in coupler frequency $\omega_{c}$ and direct coupling, i.e. $g_{13}$. We present the findings in Fig. 20.
Refer to caption(b)(c)(d)(e)(f)
Figure 20: (a)Triangle geometry of 3 qubit interaction indirectly coupled via one shared coupler. Stray couplings versus coupler C frequency of circuit (b) $ZZZ$, (c) $ZZI$, (d) $ZIZ$, (e) $IZZ$, (f) $|ZZZ|/{|ZZ|_{\max}}$. Circuit parameters are: $\omega_{1}/2\pi=4.8$ GHz, $\omega_{2}/2\pi=5.0$ GHz, $\omega_{3}/2\pi=5.1$ GHz, $\delta_{1}/2\pi=\delta_{2}/2\pi=\delta_{3}/2\pi=-330$ MHz with qubit-coupler strength $g_{1c1}/2\pi=g_{1c_{3}}/2\pi=g_{2c_{1}}/2\pi=g_{3c_{3}}/2\pi=85$ MHz and $g_{2c_{2}}/2\pi=g_{3c_{2}}/2\pi=102$ MHz, direct qubit-qubit coupling $g_{12}=g_{23}/2\pi=4$ MHz.
@@ Appendix F Three qubits and one coupler
Appendix G Cross Resonance gate simulation
We study the impact $CR$ gate on the lattice Hamiltonian of three qubits shown in Fig. 10. For this aim we use our software CirQubit to generate numerical data for the cases that experiment has not conducted, such as the static stray coupling in absence of $CR$ gate $\Omega=0$, and the case of $CR$ gate amplitude being $\Omega=31$ MHz, which is almost the double of the experimental data taken at $\Omega=16$ MHz. Results have been plotted in Fig. 21, which indicates how the Lattice Pauli coefficients of Eq. (12) depends on the frequency detuning between control and target qubits $\Delta_{st}$. Black dashed lines indicates static divergences that exist in the circuit in the static undriven circuit. In contrast, the blue dashed lines are new divergences that appear in the microwave-activated part.
Figure 21: Pauli $CR$ coefficients at three different driving amplitude: $\Omega=0$ (black), $\Omega=18$ MHz (yellow) and $\Omega=31$ MHz (pink), respectively.